# Asynchronous Microphone Array Calibration using Hybrid TDOA Information

### 12 Mar 2024 v2

@@ IV-1 Calibration Scenario
The real-world calibration scenario is shown in Fig. 4. The robot (TurtleBot3) carrying a speaker moves around a given plane trajectory whose space is $1.6m\times 2m\times 1m$. When the robot reaches the marked point, the speaker sends out a calibration signal (chirp), and there are 14 sound event locations. On the robot, the speaker is installed on a rotatable pole to change the height of the sound source. Both TDOA-S and TDOA-M are obtained by the GCC-PHAT method [14] and odometry measurements are obtained by an efficient Monocular Visual-Inertial State Estimator (VINS-Mono) [18]. There are three microphone arrays inside the trajectory, each array uses IFLYTEK M160C, which is a circular microphone array with six microphones.
@@
IV-2 Setup
We randomly set five microphone position configurations and each one is repeated three times. A certain number of microphones are selected from the three arrays randomly to form a microphone array. The advantage of extracting microphones from multiple arrays to form an array is that it can more conveniently generate a large amount of real data, i.e. thousands of data samples, and enhance experimental persuasiveness.
We conduct verification of our method in three sub-experiments corresponding to the simulation of Part A, Part B, and Part C. The real-world experiment settings are the same as that shown in Table I concerning the three parts, except that there is only one sound source trajectory with 14 sound events and TDOA noises of real-world data need to be divided based on their estimation results. Also, in “True $\mathbf{x}_{mic}$”, “random” means microphones are randomly selected from three microphone arrays and $\sigma_{tdoa}=10^{-4}s$.
IV-3 TDOA Noises Evaluation
It’s necessary to estimate the noises of TDOA-S and TDOA-M before conducting the real experiment. Because the true values of both microphone and sound locations are known, the estimated noise standard deviation of TDOA-S ($\tilde{\sigma}^{S}_{tdoa}$) and TDOA-M ($\tilde{\sigma}^{M}_{tdoa}$) are obtained based on MLE in Appendix B.
IV-A Real-World Experiment Results
In Part A and B, to ensure fairness, we select data satisfying $|\tilde{\sigma}^{S}_{tdoa}-\tilde{\sigma}^{M}_{tdoa}|<10^{-5}s$. In Part C, data is divided into five cases with different estimated TDOA noises: $\tilde{\sigma}^{S}_{tdoa},\tilde{\sigma}^{M}_{tdoa}<10^{-4}s$ (Case A), $10^{-4}s<\tilde{\sigma}^{S}_{tdoa},\tilde{\sigma}^{M}_{tdoa}<1.5\times 10^{-4}s$ (Case B), $1.5\times 10^{-4}s<\tilde{\sigma}^{S}_{tdoa},\tilde{\sigma}^{M}_{tdoa}<5\times 10^{-4}s$ (Case C), $|\tilde{\sigma}^{S}_{tdoa}-\tilde{\sigma}^{M}_{tdoa}|<10^{-5}s$ (Case D) and all TDOA-S and TDOA-M without any conditions (Case E).
[…]
Figure 5: Box plot of results in the real-world experiment: microphone location estimation errors under various microphone numbers (a), initial values noises (b), and five cases of TDOA noises (c).
V CONCLUSIONS
In the scenario of a robot carrying a sound source moving around a microphone array, this paper constructs a Graph SLAM utilizing hybrid TDOA and odometer information to simultaneously estimate microphone parameters (including microphone positions, time offsets, and clock offset rates) and sound source position. Hybrid TDOA is composed of TDOA-M and TDOA-S and the latter is inspired by sound source localization to locate microphones. TDOA-S is efficient and simple, which eliminates time offsets without generating new parameters. Both simulation and real-world experiment results consistently prove that our method is independent of the number of microphones, has low sensitivity to initial values, and has higher accuracy and robustness under various TDOA noises. In addition, simulations show that our method has a lower CRLB for microphone parameters, which explains the advantages of our method from an information theory perspective.
VI APPENDIX
VI-A Affine Transformation from $\mathbf{x}_{mic}^{S}$ to $\mathbf{x}_{mic}^{M}$
@@
$\mathbf{x}_{mic}=[\mathbf{x}_{1},\mathbf{x}_{2},\tau_{2,1},\delta_{2,1},...,\mathbf{x}_{N},\tau_{N,1},\delta_{N,1}]^{T},$ (17)
Given the definitions of two coordinate systems: $Sound$ frame and $Mic.$ frame, and the definitions of two parameter vectors $\mathbf{x}_{mic}^{S}$ and $\mathbf{x}_{mic}^{M}$, the details of this linear transformation relationship are as follows,
@@
where $\mathbf{R}$ and $\mathbf{t}$ are the rotation matrix and translation vector respectively and transfer $\mathbf{x}_{i}$ in $Sound$ frame into $Mic.$ frame. The construction of $\mathbf{A}^{M}_{S}$ and $\mathbf{b}^{M}_{S}$ are based on (1), (17) and (18).
VI-B Estimating Standard Deviation of TDOA Noise
VI-B1 Computation of $\tilde{\sigma}^{S}_{tdoa}$
Given $t^{S}_{i,j}$, $\mathbf{x}_{i}$ and $\mathbf{s}_{j}$, $i=1,2,...,N$ and $j=1,2,...,K-1$. $\tilde{t}^{S}_{i,j}$ is shown below,
@@ VI-B1 Computation of $\tilde{\sigma}^{S}_{tdoa}$
Unbiased estimation based on MLE for $\delta_{i}$ is below,
@@
VI-B2 Computation of $\tilde{\sigma}^{M}_{tdoa}$
Given $t^{M}_{i,j}$, $\mathbf{x}_{i}$ and $\mathbf{s}_{j}$, $i=2,3,...,N$ and $j=1,2,...,K$. $\tilde{t}^{M}_{i,j}$ is shown below,
@@ VI-B2 Computation of $\tilde{\sigma}^{M}_{tdoa}$
Unbiased estimation based on MLE for $\tau_{i,1},\delta_{i,1}$ are below,

### 19 Mar 2024 v3

@@ IV-1 Calibration Scenario
The real-world calibration scenario is shown in Fig. 4. The robot (TurtleBot3) carrying a speaker moves around a given plane trajectory whose space is $1.6m\times 2m\times 1m$. When the robot reaches the marked point, the speaker sends out a calibration signal (chirp), and there are 14 sound event locations. On the robot, the speaker is installed on a rotatable pole to change the height of the sound source. Both TDOA-S and TDOA-M are obtained by the GCC-PHAT method [19] and odometry measurements are obtained by an efficient Monocular Visual-Inertial State Estimator (VINS-Mono) [22]. There are three microphone arrays inside the trajectory, each array uses IFLYTEK M160C, a circular array with six microphones.
IV-2 Setup
We randomly set five microphone position configurations and each one is repeated three times. A certain number of microphones are selected from the three arrays randomly to form a microphone array. The advantage of extracting microphones from multiple arrays to form an array is that it can generate a large amount of real data more conveniently. We conduct similar comparisons with Section III (Part A, Part B, and Part C). The real-world experiment settings are the same as shown in Table I, except that there is only one sound source trajectory with 14 sound events, and TDOA noises of real-world data need to be derived based on the real data. Also, in “True $\mathbf{x}_{mic}$”, “random” means microphones are randomly selected from three microphone arrays. $\sigma_{tdoa}$ is set to $10^{-4}s$.
IV-3 TDOA Noises Evaluation
It’s necessary to estimate the noises of TDOA-S and TDOA-M before conducting the real experiment. Because the true values of both microphone and sound locations are known, the estimated noise standard deviation of TDOA-S ($\tilde{\sigma}^{S}_{tdoa}$) and TDOA-M ($\tilde{\sigma}^{M}_{tdoa}$) are obtained based on MLE in Appendix B. In Part A and B, to ensure fairness, we select data satisfying $|\tilde{\sigma}^{S}_{tdoa}-\tilde{\sigma}^{M}_{tdoa}|<10^{-5}s$. In Part C, data is divided into five cases with different estimated TDOA noises: $\tilde{\sigma}^{S}_{tdoa},\tilde{\sigma}^{M}_{tdoa}<10^{-4}s$ (Case A), $10^{-4}s<\tilde{\sigma}^{S}_{tdoa},\tilde{\sigma}^{M}_{tdoa}<1.5\times 10^{-4}s$ (Case B), $1.5\times 10^{-4}s<\tilde{\sigma}^{S}_{tdoa},\tilde{\sigma}^{M}_{tdoa}<5\times 10^{-4}s$ (Case C), $|\tilde{\sigma}^{S}_{tdoa}-\tilde{\sigma}^{M}_{tdoa}|<10^{-5}s$ (Case D) and all TDOA-S and TDOA-M without any conditions (Case E).
IV-4 Results
[…]
@@ IV-4 Results
Figure 5: Box plot of results in the real-world experiment: microphone location estimation errors under various microphone numbers (a), i.e. 4, 6, 8, and 10, initial values noise SDs (b), i.e. 0m, 0.5m, 1m, and 2m, and five cases of TDOA noises (c).
V DISCUSSION
The simulation and experimental results show that our calibration method, incorporating both TDOA-S and TDOA-M measurements, has several advantages compared to the case using only TDOA-M information. A major reason for the above outcome could be that TDOA-S actually has fewer timing parameters than TDOA-M, which helps to reduce the nonlinearity of non-linear LS and sensitivity to initial values. For lower CRLB, from an information theory perspective, compared to [15], we have more measurement information, which results in lower CRLB.
VI CONCLUSIONS
This paper is concerned with asynchronous microphone array calibration using batch SLAM. More specifically, we have introduced a new type of measurement, i.e. TDOA-S, for the above calibration problem. We have presented a simple procedure to extract TDOA—S measurements from raw audio information. We proposed to use hybrid TDOA information (both TDOA-S and TDOA-M) for calibrating asynchronous microphone arrays. Both simulation and experiment results show that our proposed method leads to more desirable calibration results than the case using only TDOA-M information. In particular, the proposed method is more independent of the number of microphones, has lower sensitivity to initial values, and has higher accuracy and robustness under various TDOA noises. The focus of our current and future work is to explore the relationship between calibration observability (eigenvalues of the Fisher information matrix) and sound source trajectory configuration and generalize the ideas in this paper to the problem of calibrating multiple microphone arrays.
VII APPENDIX
VII-A Affine Transformation from $\mathbf{x}_{mic}^{S}$ to $\mathbf{x}_{mic}^{M}$
@@
$\mathbf{x}_{mic}=[\mathbf{x}_{1},\mathbf{x}_{2},\tau_{2,1},\delta_{2,1},...,\mathbf{x}_{N},\tau_{N,1},\delta_{N,1}]^{T}.$ (17)
Given the definitions of the vectors $\mathbf{x}_{mic}^{S}$ and $\mathbf{x}_{mic}^{M}$, the details of this linear transformation relationship are as follows
@@
where $\mathbf{R}$ and $\mathbf{t}$ are the rotation matrix and translation vector respectively and transfer $\mathbf{x}_{i}$ from $Sound$ frame into $Mic.$ frame. The construction of $\mathbf{A}^{M}_{S}$ and $\mathbf{b}^{M}_{S}$ are based on (1), (17) and (18).
VII-B Estimating Standard Deviation of TDOA Noise
VII-B1 Computation of $\tilde{\sigma}^{S}_{tdoa}$
Given $t^{S}_{i,j}$, $\mathbf{x}_{i}$ and $\mathbf{s}_{j}$, $i=1,2,...,N$ and $j=1,2,...,K-1$. $\tilde{t}^{S}_{i,j}$ is shown below
@@ VII-B1 Computation of $\tilde{\sigma}^{S}_{tdoa}$
Unbiased estimation based on MLE for $\delta_{i}$ is below
@@
VII-B2 Computation of $\tilde{\sigma}^{M}_{tdoa}$
Given $t^{M}_{i,j}$, $\mathbf{x}_{i}$ and $\mathbf{s}_{j}$, $i=2,3,...,N$ and $j=1,2,...,K$. $\tilde{t}^{M}_{i,j}$ is shown below
@@ VII-B2 Computation of $\tilde{\sigma}^{M}_{tdoa}$
Unbiased estimation based on MLE for $\tau_{i,1},\delta_{i,1}$ are below